# Information Science Principles of Machine Learning: A Causal Chain Meta-Framework Based on Formalized Information Mapping

### v11

@@ C.3 Formal Representations of Information at Each Stage of Model M
$Io\left(\varphi o_{i}\right)\trianglerighteq\psi o_{i}(1\leq i\leq 11)$ (C.28)
Here, the information $Io$ enables the mapping from the initial ontological state of model M to its carrier state, that is, the initial state of M. In practice, this mapping is a complex process involving hardware implementation, software development, and hardware-software integration, inevitably requiring consumption of time and energy resources, and subject to various sources of noise. However, to simplify the discussion, we temporarily disregard these factors and explain the initial state of M using a completely ideal, noise-free, and recoverable information model. The subsequent discussions on other types of information will follow this approach and will not be repeated.
@@
$It=\left\langle ot,{Tt}_{h},S_{ot},Mt,{Tt}_{m},S_{Mt}\right\rangle$ (C.29)
Here, the ontology $ot$ refers to the object described by the training information, and the occurrence time ${Tt}_{h}$ denotes the point at which this object attains its corresponding state attributes. The ontological state $S_{ot}(ot,{Tt}_{h})$ represents the attributes of the object at time ${Tt}_{h}$, which are described by a set of well-formed formulas:
@@
The carrier $Mt$ serves as the interface device through which model M receives training data, and ${Tt}_{m}$ is the time at which model M is trained. The carrier state $S_{Mt}(Mt,{Tt}_{m})$ is then described by the set of well-formed formulas:
$It\left(\varphi t_{i}\right)\trianglerighteq\psi t_{i}(1\leq i\leq 2)$ (C.33)
@@
is a one-to-one enabling mapping from $S_{ou}\left(ou,{Tu}_{h}\right)$ onto M, representing the state of M after completing its training.
@@
The carrier $Mq$ for $Iq$ serves as the input device for M, which may or may not be the same as the carrier $Mt$ for the training information $It$. The state $S_{Mq}(Mq,{Tq}_{m})$ is a one-to-one enabling mapping from $S_{oq}(oq,{Tq}_{h})$ onto $Mq$.
@@
In this case, the output information $Ir=\left\langle or,{Tr}_{h},S_{or},Mr,{Tr}_{m},S_{Mr}\right\rangle$ has an ontological state defined as:
@@
The carrier state $S_{Mr}(Mr,{Tr}_{m})$ is a one-to-one enabling mapping from $S_{or}\left(or,{Tr}_{h}\right)$ to the output device of M.
@@
is a one-to-one enabling mapping from $S_{ov}\left(ov,{Tv}_{h}\right)$ onto M.
At this point, based on MLT-MF, we have constructed the formal well-formed formula representations for all ontological and carrier states of each component of the feedforward neural network model M, as illustrated in Fig. 7. Although M is a simple model, it contains all the essential elements of a neural network, and the discussion has not relied on any special-case treatment of M. This demonstrates that MLT-MF is capable of supporting the construction of well-formed, formal representations for a wide variety of neural networks, providing a solid foundation for detailed analysis and optimization.

### v12

@@ C.3 Formal Representations of Information at Each Stage of Model M
$Io:\varphi o_{i}\rightrightarrows\psi o_{i}\quad(1\leq i\leq 11)$ (C.28)
Here, the information $Io$ enables the mapping from the initial ontological state of model M to its carrier state, that is, the initial state of M. In practice, this mapping is a complex process involving hardware implementation, software development, and hardware-software integration, inevitably requiring consumption of time and energy resources, and subject to various sources of noise. However, to simplify the discussion, we temporarily disregard these factors and explain the initial state of M using a completely ideal, and noise-free information model. The subsequent discussions on other types of information will follow this approach and will not be repeated.
@@
$It:S_{ot}(ot,Tt_{h})\rightrightarrows S_{Mt}(Mt,Tt_{m})$ (C.29)
Here, the ontology $ot$ refers to the object described by the training information, and the occurrence time ${Tt}_{h}$ denotes the point at which this object attains its corresponding state attributes. The ontological state $S_{ot}(ot,{Tt}_{h})$ represents the attributes of the object at time ${Tt}_{h}$, which are described by a set of WFFs:
@@
The carrier $Mt$ serves as the interface device through which model M receives training data, and ${Tt}_{m}$ is the time at which model M is trained. The carrier state $S_{Mt}(Mt,{Tt}_{m})$ is then described by the set of WFFs:
$It:\varphi t_{i}\rightrightarrows\psi t_{i}\quad(1\leq i\leq 2)$ (C.33)
@@
is a noisy-free enabling mapping from $S_{ou}\left(ou,{Tu}_{h}\right)$ onto M, representing the state of M after completing its training.
@@
The carrier $Mq$ for $Iq$ serves as the input device for M, which may or may not be the same as the carrier $Mt$ for the training information $It$. The state $S_{Mq}(Mq,{Tq}_{m})$ is a noisy-free enabling mapping from $S_{oq}(oq,{Tq}_{h})$ onto $Mq$.
@@
In this case, the output information $Ir:S_{or}(or,Tr_{h})\rightrightarrows S_{Mr}(Mr,Tr_{m})$ has an ontological state defined as:
@@
The carrier state $S_{Mr}(Mr,{Tr}_{m})$ is a noisy-free enabling mapping from $S_{or}\left(or,{Tr}_{h}\right)$ to the output device of M.
@@
is a noisy-free enabling mapping from $S_{ov}\left(ov,{Tv}_{h}\right)$ onto M.
At this point, based on MLT-MF, we have constructed the formal well-formed formula representations for all ontological and carrier states of each component of the feedforward neural network model M, as illustrated in Fig. 7. Although M is a simple model, it contains all the essential elements of a neural network, and the discussion has not relied on any special-case treatment of M. This demonstrates that MLT-MF is capable of supporting the construction of WFF representations for a wide variety of neural networks, providing a solid foundation for detailed analysis and optimization.